Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Hard Manhattan Problem

Expert replies
by kakz » Sun Feb 26, 2012 1:40 am
Two integers between 1 and 100, inclusive, each randomly and independently chosen, are either added or multiplied, with an equal chance of either operation. What is the probability that the result is even?
(A)1/3
(B)1/2
(C)5/8
(D)2/3
(E)7/8
Join the discussion
Source: — Problem Solving |

by piyush272 » Sun Feb 26, 2012 3:46 am
its easier than you think.!
C) 5/8

total cases when two nos. are randomly picked:
if both are>> odd,odd=sum> EVEN
product> ODD
odd,even=sum> ODD
product> EVEN
even,even=sum> EVEN
product> EVEN
even,odd=sum> ODD
product> EVEN
hence favorable cases= 5
total cases= 8
probability = 5/8

all the best.!
Last edited by piyush272 on Sun Feb 26, 2012 5:46 am, edited 1 time in total.
Join the discussion

by kakz » Sun Feb 26, 2012 4:30 am
That is not correct OA is C
Join the discussion

by GMATGuruNY » Sun Feb 26, 2012 4:43 am
kakz wrote:Two integers between 1 and 100, inclusive, each randomly and independently chosen, are either added or multiplied, with an equal chance of either operation. What is the probability that the result is even?
(A)1/3
(B)1/2
(C)5/8
(D)2/3
(E)7/8
P(even) = 1 - P(odd).

The following are ways to get an ODD result.

Case 1: odd*odd = odd.
P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8.

Case 2: odd+even = odd OR even+odd = odd.
P(odd number and addition and even number) = 1/2 * 1/2 * 1/2 = 1/8.
P(even number and addition and odd number) = 1/2 * 1/2 * 1/2 = 1/8.
1/8 + 1/8 = 2/8.

P(even) = 1 - (1/8 + 2/8) = 5/8.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Mr.Hollywood » Mon Feb 27, 2012 8:52 pm
GMATGuruNY wrote:
kakz wrote:Two integers between 1 and 100, inclusive, each randomly and independently chosen, are either added or multiplied, with an equal chance of either operation. What is the probability that the result is even?
(A)1/3
(B)1/2
(C)5/8
(D)2/3
(E)7/8
P(even) = 1 - P(odd).

The following are ways to get an ODD result.

Case 1: odd*odd = odd.
P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8.

Case 2: odd+even = odd OR even+odd = odd.
P(odd number and addition and even number) = 1/2 * 1/2 * 1/2 = 1/8.
P(even number and addition and odd number) = 1/2 * 1/2 * 1/2 = 1/8.
1/8 + 1/8 = 2/8.

P(even) = 1 - (1/8 + 2/8) = 5/8.

The correct answer is C.
Hi forgive my stupidity. Could you further explain the part where: "P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8" where did the third 1/2 come from?

Thank you!
Join the discussion

by GMATGuruNY » Mon Feb 27, 2012 9:06 pm
Mr.Hollywood wrote:
GMATGuruNY wrote:
kakz wrote:Two integers between 1 and 100, inclusive, each randomly and independently chosen, are either added or multiplied, with an equal chance of either operation. What is the probability that the result is even?
(A)1/3
(B)1/2
(C)5/8
(D)2/3
(E)7/8
P(even) = 1 - P(odd).

The following are ways to get an ODD result.

Case 1: odd*odd = odd.
P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8.

Case 2: odd+even = odd OR even+odd = odd.
P(odd number and addition and even number) = 1/2 * 1/2 * 1/2 = 1/8.
P(even number and addition and odd number) = 1/2 * 1/2 * 1/2 = 1/8.
1/8 + 1/8 = 2/8.

P(even) = 1 - (1/8 + 2/8) = 5/8.

The correct answer is C.
Hi forgive my stupidity. Could you further explain the part where: "P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8" where did the third 1/2 come from?

Thank you!
Case 1: odd*odd = odd.
P(first number is odd) = 1/2. (Since 1/2 of the integers between 1 and 100, inclusive, are odd.)
P(multiplication) = 1/2. (Since the integers are either multiplied or added, with an equal chance of either operation.)
P(second number is odd) = 1/2.
Since we want all of these events to happen together, we multiply the fractions:
1/2 * 1/2 * 1/2 = 1/8.

The same reasoning was applied to the other cases.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by krusta80 » Tue Feb 28, 2012 4:06 pm
1. With multiplication, we get an even result whenever at least one of the numbers is even.

Let A denote the first number being even
Let B denote the second being even

P(A or B) = P(A) + P(B) - P(A and B) = 1/2 + 1/2 - 1/4 = 3/4


2. With addition, we get an even result whenever both numbers are even OR both numbers are odd.

Keeping with the same notation above...

P(A and B) = 1/4
P(not A and not B) = 1/4

Since the above possibilities are mutually exclusive, we simply add them together to get the probability of one or the other occurring.

Finally, we average the two cases together, since multiplication or addition can occur with equal chance: (3/4 + 1/2) / 2 = 5/8.
Join the discussion

by Kyan » Fri Dec 21, 2012 7:38 pm
Mr.Hollywood wrote:
GMATGuruNY wrote:
kakz wrote:Two integers between 1 and 100, inclusive, each randomly and independently chosen, are either added or multiplied, with an equal chance of either operation. What is the probability that the result is even?
(A)1/3
(B)1/2
(C)5/8
(D)2/3
(E)7/8
P(even) = 1 - P(odd).

The following are ways to get an ODD result.

Case 1: odd*odd = odd.
P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8.

Case 2: odd+even = odd OR even+odd = odd.
P(odd number and addition and even number) = 1/2 * 1/2 * 1/2 = 1/8.
P(even number and addition and odd number) = 1/2 * 1/2 * 1/2 = 1/8.
1/8 + 1/8 = 2/8.

P(even) = 1 - (1/8 + 2/8) = 5/8.

The correct answer is C.
Hi forgive my stupidity. Could you further explain the part where: "P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8" where did the third 1/2 come from?

Thank you!
I came across this question today and just stumbled upon this great explanation. I just have one quick question. How can we say that the probability of selecting the second odd number is also 1/2? Assuming we have 50 odd numbers in a group of 100 integers, shouldn't the probability of randomly selecting the second odd number be 49/99? Instead of 1/2?

Thanks!
Join the discussion

by GMATGuruNY » Fri Dec 21, 2012 9:15 pm
Kyan wrote:
Mr.Hollywood wrote:
GMATGuruNY wrote:
kakz wrote:Two integers between 1 and 100, inclusive, each randomly and independently chosen, are either added or multiplied, with an equal chance of either operation. What is the probability that the result is even?
(A)1/3
(B)1/2
(C)5/8
(D)2/3
(E)7/8
P(even) = 1 - P(odd).

The following are ways to get an ODD result.

Case 1: odd*odd = odd.
P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8.

Case 2: odd+even = odd OR even+odd = odd.
P(odd number and addition and even number) = 1/2 * 1/2 * 1/2 = 1/8.
P(even number and addition and odd number) = 1/2 * 1/2 * 1/2 = 1/8.
1/8 + 1/8 = 2/8.

P(even) = 1 - (1/8 + 2/8) = 5/8.

The correct answer is C.
Hi forgive my stupidity. Could you further explain the part where: "P(odd number and multiplication and odd number) = 1/2 * 1/2 * 1/2 = 1/8" where did the third 1/2 come from?

Thank you!
I came across this question today and just stumbled upon this great explanation. I just have one quick question. How can we say that the probability of selecting the second odd number is also 1/2? Assuming we have 50 odd numbers in a group of 100 integers, shouldn't the probability of randomly selecting the second odd number be 49/99? Instead of 1/2?

Thanks!
The two integers are INDEPENDENTLY chosen.
This means that neither selection restricts the other.
The first number selected can be any integer between 1 and 100, inclusive, as can the second number.
Thus -- in each case -- the probability of selecting an odd integer = 1/2.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion